An environmental economic consulting firm is hired to measure the negative externalities associated with the pollution from an industry. The consultants calculate the marginal social cost of production to be MSC 20+30 and the marginal private cost of production to be MPC Q+30. The market demand curve can be expressed as P 60-Q. If the consultants have accurately measured the impact of the pollution externality, the deadweight loss from producing at the market equilibrium is A. $30 ов B. $37.5. OC. $55 D. $67.5.

Step by step explanations are appreciated. I did MSC-MPC but I just got Q.
Answer #1

Option (B).

Social outcome is obtained by equating Demand and MSC.

60 - Q = 2Q + 30

3Q = 30

Q = 10

P = 60 - 10 = $50

Market outcome is obtained by equating Demand and MPC.

60 - Q = Q + 30

2Q = 30

Q = 15

When Q = 15,

MSC = (2 x 15) + 30 = 30 + 30 = $60

MPC = 15 + 30 = $45

Deadweight loss = (1/2) x (MSC - MPC) x Change in output = (1/2) x $(60 - 45) x (15 - 10) = (1/2) x $15 x 5 = $37.5